OpenStudy (anonymous):

Quick Trig Question: Is square root (1 - cos^x) = -sin x, true? If so which quadrants does it lie in? If it's true I know it lies in III and IV, am I right?

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @mathmale

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm \sqrt{1-\cos^2x}=-\sin x\] Let's simplify the left side,\[\Large\rm \sqrt{\sin^2x}=-\sin x\]and further,\[\Large\rm \pm \sin x=-\sin x\]So we can see that this is `sometimes` true, yes? And yes, you are correct, sin x is negative in quadrants 3 and 4. So that's specifically where it's true.

OpenStudy (anonymous):

Okay, the choices are this: False True; quadrants I & IV True; quadrants II & III True; quadrants III & IV

OpenStudy (anonymous):

On Yahoo Answers someone said tht "True, quadrants I & IV" is correct?

OpenStudy (anonymous):

WAIT nevermind it was a different question.

OpenStudy (anonymous):

But as for my answer, the fourth choice would be correct?

zepdrix (zepdrix):

Yes looks correct :) Here is a graph to help you verify. https://www.desmos.com/calculator/rjcyol5znp I graphed the left and right side separately. You can hover over certain points and see that they start to overlap at pi, and then stop overlapping at 2pi.

OpenStudy (anonymous):

Awesome, thank you so much again! :)